# Conceptualization, Data Curation, Software, Visualization, Writing - original draft

### 01 Mar 2026 v5

@@ Abstract
Dynamic line rating (DLR) enables greater utilization of existing transmission lines by leveraging real-time weather data. However, the elevated temperature operation (ETO) of conductors under DLR is often overlooked, despite its long-term impact on conductor health. This paper addresses this issue by 1) quantifying risk-based depreciation costs associated with ETO and 2) proposing a Conductor Health-Aware Unit Commitment (CHA-UC) that internalizes these costs in operational decisions. CHA-UC incorporates a robust linear approximation of conductor temperature and integration of expected depreciation costs due to hourly ETO into the objective function. Case studies on the Texas 123-bus backbone test system using NOAA weather data demonstrate that the proposed CHA-UC model reduces the total cost by 0.74% and renewable curtailment by 85% compared to static line rating (SLR) and outperforms quantile regression forest-based methods, while conventional DLR operation without risk consideration resulted in higher costs due to excessive ETO. Further analysis of the commitment decisions and the line temperature statistics confirms that the CHA-UC achieves safer line flows by shifting generator commitments. Finally, we examine the emergent correlation behaviors arising between wind generation and DLR forecast errors, and show that CHA-UC adaptively manages this effect by relaxing flows for risk-hedging conditions while tightening flows for risk-amplifying ones.
keywords:
[…]
@@
$b\in\mathcal{B}$ Set of buses. $l\in\mathcal{L}$ Set of all transmission lines where $\mathcal{L}^{\text{DLR/SLR}}$ denote sets of DLR/SLR applied lines $i\in\mathcal{I}$ Set of thermal generating units. $o\in\mathcal{O}_{i}$ Set of startup segments for unit $i$. $w\in\mathcal{W}_{b}$ Set of wind farms connected at bus $b$. $t\in\mathcal{T}$ Set of dispatch time periods. $s\in\mathcal{S}$ Set of uncertainty scenarios. s(l),r(l)s(l),\,r(l) Sending‐ and receiving‐end buses of line $l$.
@@
$A_{l,t}$, $B_{l,t}$ Slope and intercept of linear fit of loading-temperature for line $l$ [°C/MW, °C]. $\hat{A}_{l,t}$, $\hat{B}_{l,t}$ DA forecast of slope and intercept of the linear fit for line $l$ at time $t$ [°C/MW, °C]. $C^{\rm Dep}_{l,t}(\cdot)$ Line depreciation cost function [\$/°C]. $C^{\mathrm{P}}_{i}(\cdot)$ Quadratic fuel cost function [\$/MWh]. $C^{\mathrm{r}+}_{i},\,C^{\mathrm{r}-}_{i}$ Up and down‐reserve costs [\$/MWh]. $C^{\mathrm{SU}}_{i,o},\,C^{\mathrm{SD}}_{i}$ Startup and shutdown costs [\$/MWh]. $C^{\rm VoLL}$ Value of lost load [\$/MWh]. $\underline{F}^{\mathrm{SLR}}_{l},\,\overline{F}^{\mathrm{SLR}}_{l}$ SLR of line $l$ [MW]. $\underline{F}^{\mathrm{DLR}}_{l,t},\,\overline{F}^{\mathrm{DLR}}_{l,t}$ RT DLR of line $l$, time $t$ [MW]. $\hat{\underline{F}}^{\mathrm{DLR}}_{l,t},\,\hat{\overline{F}}^{\mathrm{DLR}}_{l,t}$ DA DLR forecasts of line $l$, time $t$ [MW]. $\underline{G}_{i},\,\overline{G}_{i}$ Minimum, maximum power output [MW]. $P^{\rm d}_{b,t}$ Load at bus $b$ and time $t$ [MW]. $P^{\rm wind}_{w,t}$ RT WP generation [MW]. $\hat{P}^{\rm wind}_{w,t}$ DA Forecast WP generation [MW]. $\Pr_{s}$ Probability of scenario $s$. $R^{\text{min},+}_{t}$ Required system total up reserve. $RU_{i},\,RD_{i}$ Ramp‐up and ramp‐down limits [MW/h]. $SU_{i},\,SD_{i}$ Startup and shutdown capability [MW]. $T^{\mathrm{SU}}_{i,o}$ Hours defining the segment $o$ limits [h]. $TU_{i},\,TD_{i}$ Minimum up‐ and down‐times of unit $i$ [h]. $X_{l}$ Reactance of line $l$ [p.u.]. $\underline{\Theta}_{l},\,\overline{\Theta}_{l}$ Phase‐angle difference limits on line $l$ [rad]. $\xi_{l,t,s}$ DLR forecast error scenario $s$ [MW]. $\omega_{w,t,s}$ WP forecast error scenario $s$ [MW]. $K_{l,t}$ DLR security margin multiplier.
@@
$g_{i,t}$ DA dispatch of unit $i$ at time $t$ [MW]. $\hat{g}^{\rm adj}_{i,t,s}$,   $g^{\rm adj}_{i,t}$ DA and RT adjusted generation of unit $i$ at time $t$ [MW]. $p_{l,t}$ DA power flow on line $l$ at time $t$ [MW]. $\hat{p}^{\rm adj}_{l,t,s}$, $p^{\rm adj}_{l,t}$ DA and RT adjusted power flow of line $l$ at time $t$ [MW]. $p^{\rm cur}_{w,t}$ DA wind generation curtailment [MW]. $\hat{p}^{\rm cur,adj}_{w,t,s}$, $p^{\rm cur,adj}_{w,t}$ DA and RT adjusted wind generation curtailment at farm $w$, time $t$ [MW]. $\hat{p}^{\rm shed}_{b,t}$ DA load shedding at bus $b$, time $t$ [MW]. $\hat{p}^{\rm shed,adj}_{b,t,s}$, $p^{\rm shed,adj}_{b,t}$ DA and RT adjusted load shedding at bus $b$, time $t$ [MW]. $r^{\text{avail},+}_{i,t}$ DA available online up reserve [MW]. $\hat{r}^{+}_{i,t,s},\hat{r}^{-}_{i,t,s}$ DA up and down‐reserve activation given scenario $s$ for unit $i$, time $t$, [MW]. $r^{+}_{i,t},r^{-}_{i,t}$ RT up and down‐reserve activation for unit $i$, time $t$, [MW]. $u_{i,t}$ Binary commitment variable $u_{i,t}^{\rm SU}$$u_{i,t}^{\rm SD}$ Binary startup/shutdown variable. $\delta_{i,o,t}$ Activation of startup‐segment $o$. $\theta_{b,t}$ DA voltage angle [rad]. $\hat{\theta}^{\rm adj}_{b,t,s}$$\theta^{\rm adj}_{b,t}$ DA and RT adjusted voltage angle [rad]. $\hat{\tau}_{l,t,s}$ DA conductor temperature given scenario $s$ on line $l$, time $t$ [°C]. $\tau_{l,t}$ RT temperature of conductor [°C].
1 Introduction
@@
The large-scale integration of renewable energy sources (RES) into power systems remains a major technical challenge, particularly due to the spatial mismatch between high-potential generation sites and demand centers. Areas with high RES potential are far from demand, and efficient RES integration plans entail intense transmission expansion Impram2020, Zhang2025. Case studies in Texas indicate that transmission expansion costs rise significantly when RES are integrated due to their remote locations Herding2024 while new transmission build-outs, in turn, significantly enhance wind project profitability Kemp2023. Moreover, transmission construction lags generation: wind and solar projects can be built within three years, whereas new transmission requires five to ten years Donoho2015, Clapin2024. This mismatch causes grid congestion, increased operation costs, and curtailed renewable generation. For instance, in 2024, 4% of WP generation and 6% of solar generation were curtailed in Texas Potomac2025, while in Korea, four out of ten RES producers could not connect to the grid, causing an estimated \$6.7 billion in losses Park2023. Meanwhile, rapidly growing electricity demand from AI data-center infrastructure heightens the urgency of transmission expansion, as transmission delays are major binding limits on timely load additions Lal2025. Given these challenges, maximizing the use of existing transmission lines becomes essential for both RES integration and emerging load growth.
@@
Dynamic line rating (DLR) has emerged as a promising solution that exploits real-time weather conditions to increase the utilization of existing transmission facilities. For overhead transmission lines under 100 km, the transmission capacity is mostly restricted by its thermal limits. Unlike static line rating (SLR), which assumes conservative worst-case weather conditions (e.g., 0.5 m/s wind speed, 40°C ambient temperature), DLR adjusts ampacity dynamically. The synergy with wind power is particularly strong, as both depend on wind speed, a spatially correlated factor Glaum2023. Authors of Wallnerstrom2015, Lee2022 and He2025 all prove that applying DLR is effective in reducing WP curtailment. Especially, Lee2022 shows that the system operator may save 776 M\$ per year from congestion costs in Texas by introducing DLR and He2025 further validates this benefit using frequency dynamic simulations. Federal Energy Regulatory Commission (FERC) is working to advance its framework on DLR. In 2024, FERC issued an Advance Notice of Proposed Rulemaking (ANOPR) related to DLR FERC2024. This ANOPR builds on Order 881 FERC2021, detailing the addition of wind and solar factors to ambient-adjusted ratings.
@@
DLR, however, poses risks of sag and accelerated conductor degradation, especially under elevated temperature operation (ETO). Moreover, the correlation between wind generation and DLR forecasts can amplify the risk. This risk was not considered in Wallnerstrom2015 and Lee2022. Prior work has attempted to address forecast uncertainty using chance-constrained OPF Bucher2013, Gaussian copulas Viafora2020, or distributionally robust formulations Wang2018. Park et al. Park2018 extended this to stochastic unit commitment, showing that commitment decisions differ under SLR and DLR. Other works proposed risk-averse forecasts using quantile or super-quantile regression Kirilenko2021. In the actual world, FERC requires a quantile forecast value with a confidence level of 98%, meaning that the actual value of DLR should be greater than the forecast value 98% of the time FERC2024. The authors of Kim2025 obtain quantile forecast values of DLR using a graph convolutional long short-term memory model.
@@
Despite these advances, existing models do not account for the actual costs of ETO, even though ETO occurs under both SLR and DLR Adomah2000. Without quantifying ETO-induced depreciation, models remain mathematically robust but economically incomplete. Previous works have handled the effect of ETO, and some have expanded it to DLR, but no works have integrated ETO costs into decision making such as OPF or UC. Harvey’s seminal work Harvey1972 modeled the effect of ETO on aluminum conductors through an empirical equation for tensile strength loss (LoTS) as a function of conductor temperature and elapsed time. This model has been widely adopted to model ETO, including Musilek et al. Musilek2012, EPRI EPRI2005 and PJM PJM2022. Musilek2012 analyzed the historical aging of a conductor using meteorological data. Moreover, Teh2019 modeled the risk of DLR using the LoTS model and conducted a sensitivity analysis. However, this research does not consider the actual operation and performs only a simple calculation based on one line. Other depreciation models, such as fretting-based reliability analysis Morozovska2020, rely on strong assumptions with limited empirical support. To the best of our knowledge, no existing optimization framework integrates realistic conductor health costs into OPF or UC formulations.
@@
Incorporating ETO into optimization requires embedding a conductor temperature estimation model. However, because external climate conditions are highly dynamic, and the current-temperature relationship is multimodal (affected by many weather components) and nonlinear, direct integration into optimization is impractical. Previous studies have proposed conductor temperature models within OPF formulations Kirilenko2021, Ngoko2018, Dai2024. While these works employ current–temperature proxies, they do not capture the consequences of ETO, leaving a key gap in the literature. In addition, most adopt day-ahead quadratic (or biquadratic) fits of the current–temperature relationship, which present three drawbacks: (i) they can become computationally intensive when combined with stochastic optimization, (ii) they require explicit absolute (or quadratic) expressions of current Kirilenko2021 and (iii) they may underestimate conductor temperature, undermining risk management Ngoko2018.
@@
This paper addresses this gap by developing a day-ahead stochastic UC model that accounts for (i) DLR operation, (ii) power system uncertainties, and (iii) the long-term costs of ETO. Incorporating ETO costs is challenging due to the state dependency of conductor depreciation models and the lack of a closed-form solution to the loading-temperature relationship. To overcome these challenges, we introduce a piecewise linear conductor depreciation cost function and a linear line loading–temperature proxy. The main contributions of this paper are as follows:
@@
    Modeling of ETO effects: We develop tractable models to capture the long-term effects of elevated temperature operation (ETO), including (i) a robust and computationally efficient line loading–temperature proxy, and (ii) a risk-base piecewise conductor depreciation cost function.
@@
    Validation on a realistic test system: We evaluate the CHA-UC on the synthetic Texas 123-bus backbone test system (TX-123BT) under multiple sources of uncertainty. Results highlight differences in commitment decisions, demonstrate improved handling of DLR risks, and show that CHA-UC effectively addresses the correlation between wind generation and DLR forecast errors.
2 Conductor cost depreciation model
This section quantifies the long-term effects of elevated conductor temperature operation (ETO) and formulates a depreciation model suitable for optimization.
Figure 1: An illustrative example of LoTS with a conductor with exposed temperatures: 105°C, 109°C, 112°C each for two hours. The path of total degradation calculation is shadowed on trajectories of different temperatures Musilek2012.
2.1 Loss of tensile strength (LoTS) for ACSR lines
Long-term ETO causes conductor degradation, measured as the loss of tensile strength (LoTS). For aluminum strands, LoTS (%) is empirically modeled as Harvey1972:
$L^{\mathrm{Al}}=100-A\,t^{-\frac{25.4}{d}(0.001\tau-0.095)},$ (1)
where $A=\min(100,\,134-0.24\tau)$, $\tau$ is conductor temperature (°C), $t$ is exposure time (h), and $d$ is conductor diameter (mm). For aluminum conductor steel-reinforced (ACSR) lines, the corresponding LoTS is approximated as Musilek2012:

### 14 Apr 2026 v6

# Integrating Conductor Health into Dynamic Line Rating and Unit Commitment under Wind Uncertainty

@@ Abstract
Dynamic line rating (DLR) enables greater utilization of existing transmission lines by leveraging real-time weather data. However, the elevated temperature operation (ETO) of conductors under DLR, particularly in the presence of uncertainty, is often overlooked, despite its long-term impact on conductor health. This paper addresses ETO under DLR and wind power uncertainty by 1) quantifying risk-based depreciation costs associated with ETO, 2) characterizing correlation-driven ETO risk from wind power and DLR forecast errors, and 3) proposing a Conductor Health-Aware Unit Commitment (CHA-UC) that internalizes these costs in operational decisions. CHA-UC incorporates a robust linear approximation of conductor temperature and integrates expected depreciation costs due to hourly ETO into the objective function. Case studies on the Texas 123-bus backbone test system demonstrate that the proposed CHA-UC model reduces the total cost by 0.75% and renewable curtailment by 82% compared to static line rating (SLR) and outperforms quantile regression forest-based methods, while conventional DLR operation without risk consideration resulted in higher costs due to excessive ETO. Further analysis shows that CHA-UC achieves safer line utilization by shifting generator commitments and endogenously adapting to uncertainty correlation, relaxing flows under risk-hedging conditions and tightening flows under risk-amplifying conditions.
@@
$b\in\mathcal{B}$
Set of buses.
$l\in\mathcal{L}$
Set of all transmission lines where $\mathcal{L}^{\text{DLR/SLR}}$ denote sets of DLR/SLR applied lines
$i\in\mathcal{I}$
Set of thermal generating units.
$o\in\mathcal{O}_{i}$
Set of startup segments for unit $i$.
$w\in\mathcal{W}_{b}$
Set of wind farms connected at bus $b$.
$t\in\mathcal{T}$
Set of dispatch time periods.
$s\in\mathcal{S}$
Set of uncertainty scenarios.
s(l),r(l)s(l),\,r(l)
Sending‐ and receiving‐end buses of line $l$.
@@
$A_{l,t}$, $B_{l,t}$
Slope and intercept of linear fit of loading-temperature for line $l$ [°C/MW, °C].
$\hat{A}_{l,t}$, $\hat{B}_{l,t}$
DA forecast of slope and intercept of the linear fit for line $l$ at time $t$ [°C/MW, °C].
$C^{\rm Dep}_{l,t}(\cdot)$
Line depreciation cost function [\$/°C].
$C^{\mathrm{P}}_{i}(\cdot)$
Quadratic fuel cost function [\$/MWh].
$C^{\mathrm{r}+}_{i},\,C^{\mathrm{r}-}_{i}$
Up and down‐reserve costs [\$/MWh].
$C^{\mathrm{SU}}_{i,o},\,C^{\mathrm{SD}}_{i}$
Startup and shutdown costs [\$/MWh].
$C^{\rm VoLL}$
Value of lost load [\$/MWh].
$\underline{F}^{\mathrm{SLR}}_{l},\,\overline{F}^{\mathrm{SLR}}_{l}$
SLR of line $l$ [MW].
$\underline{F}^{\mathrm{DLR}}_{l,t},\,\overline{F}^{\mathrm{DLR}}_{l,t}$
RT DLR of line $l$, time $t$ [MW].
$\hat{\underline{F}}^{\mathrm{DLR}}_{l,t},\,\hat{\overline{F}}^{\mathrm{DLR}}_{l,t}$
DA DLR forecasts of line $l$, time $t$ [MW].
$\underline{G}_{i},\,\overline{G}_{i}$
Minimum, maximum power output [MW].
$P^{\rm d}_{b,t}$
Load at bus $b$ and time $t$ [MW].
$P^{\rm wind}_{w,t}$
RT WP generation [MW].
$\hat{P}^{\rm wind}_{w,t}$
DA Forecast WP generation [MW].
$\Pr_{s}$
Probability of scenario $s$.
$R^{\text{min},+}_{t}$
Required system total up reserve.
$RU_{i},\,RD_{i}$
Ramp‐up and ramp‐down limits [MW/h].
$SU_{i},\,SD_{i}$
Startup and shutdown capability [MW].
$T^{\mathrm{SU}}_{i,o}$
Hours defining the segment $o$ limits [h].
$TU_{i},\,TD_{i}$
Minimum up‐ and down‐times of unit $i$ [h].
$X_{l}$
Reactance of line $l$ [p.u.].
$\underline{\Theta}_{l},\,\overline{\Theta}_{l}$
Phase‐angle difference limits on line $l$ [rad].
$\xi_{l,t,s}$
DLR forecast error scenario $s$ [MW].
$\omega_{w,t,s}$
WP forecast error scenario $s$ [MW].
$K_{l,t}$
DLR security margin multiplier.
@@
$g_{i,t}$
DA dispatch of unit $i$ at time $t$ [MW].
$\hat{g}^{\rm adj}_{i,t,s}$,   $g^{\rm adj}_{i,t}$
DA and RT adjusted generation of unit $i$ at time $t$ [MW].
$p_{l,t}$
DA power flow on line $l$ at time $t$ [MW].
$\hat{p}^{\rm adj}_{l,t,s}$, $p^{\rm adj}_{l,t}$
DA and RT adjusted power flow [MW].
$p^{\rm cur}_{w,t}$
DA wind generation curtailment [MW].
$\hat{p}^{\rm cur,adj}_{w,t,s}$, $p^{\rm cur,adj}_{w,t}$
DA and RT adjusted wind generation curtailment at farm $w$, time $t$ [MW].
$\hat{p}^{\rm shed}_{b,t}$
DA load shedding at bus $b$, time $t$ [MW].
$\hat{p}^{\rm shed,adj}_{b,t,s}$, $p^{\rm shed,adj}_{b,t}$
DA, RT adjusted load shedding [MW].
$r^{\text{avail},+}_{i,t}$
DA available online up reserve [MW].
$\hat{r}^{+}_{i,t,s},\hat{r}^{-}_{i,t,s}$
DA up and down‐reserve activation given scenario $s$ for unit $i$, time $t$, [MW].
$r^{+}_{i,t},r^{-}_{i,t}$
RT up, down‐reserve activation [MW].
$u_{i,t}$
Binary commitment variable
$u_{i,t}^{\rm SU}$$u_{i,t}^{\rm SD}$
Binary startup/shutdown variable.
$\delta_{i,o,t}$
Activation of startup‐segment $o$.
$\theta_{b,t}$
DA voltage angle [rad].
$\hat{\theta}^{\rm adj}_{b,t,s}$$\theta^{\rm adj}_{b,t}$
DA and RT adjusted voltage angle [rad].
$\hat{\tau}_{l,t,s}$
DA conductor temperature given scenario $s$ on line $l$, time $t$ [°C].
$\tau_{l,t}$
RT temperature of conductor [°C].
I Introduction
@@
The large-scale integration of renewable energy sources (RES) into power systems remains a major technical challenge, particularly due to the spatial mismatch between high-potential generation sites and demand centers. Areas with high RES potential are far from demand, and efficient RES integration plans entail intense transmission expansion [Impram2020]. Case studies in Texas indicate that transmission expansion costs rise significantly when RES are integrated due to their remote locations [Herding2024]. Moreover, transmission construction lags generation construction: wind and solar projects can be built within three years, whereas new transmission requires five to ten years [Donoho2015, Clapin2024]. This mismatch causes increased operation costs, renewable generation curtailment, and grid congestion. Given these challenges, maximizing the use of existing transmission lines is essential for RES integration.
@@
Dynamic line rating (DLR) has emerged as a promising solution that exploits real-time weather conditions to increase the utilization of existing transmission facilities. Unlike static line rating (SLR), which assumes conservative worst-case weather conditions (e.g., 0.5 m/s wind speed, 40°C ambient temperature), DLR adjusts ampacity dynamically. The synergy with wind power is particularly strong, as both depend on wind speed, a spatially correlated factor [Glaum2023]. Authors of [Wallnerstrom2015] and [Lee2022] both prove that applying DLR is effective in reducing WP curtailment. Especially, [Lee2022] shows that the system operator may save 776 M\$ per year from congestion costs in Texas by introducing DLR. In 2024, FERC issued an Advance Notice of Proposed Rulemaking on DLR [FERC2024], building on Order 881 [FERC2021], which expanded line ratings beyond ambient-adjusted approaches. These actions foreshadow the expanding implementation and importance of DLR.
@@
However, DLR can increase the risk of accelerated conductor degradation, particularly under elevated temperature operation (ETO) and in corrosive environments. Thermal annealing and corrosion are identified as key factors limiting conductor life; corrosive environments can shorten conductor life by a factor of 1.17 [IEC2015]. Industrial conductor failure cases[AusNet2025] further highlight the need to account for the risk of ETO and corrosion in grid operation. Moreover, while DLR exhibits strong synergy with wind power, the correlation between WP and DLR forecasts can act as a source of risk amplification, not merely operational benefit. However, this risk was not considered in [Wallnerstrom2015] and [Lee2022]. Prior work has attempted to address forecast uncertainty using copula-based chance-constraints [Viafora2020] or distributionally robust formulations [Wang2018]. Park et al. [Park2018] extended this to stochastic unit commitment, showing that commitment decisions differ under SLR and DLR. Other works proposed risk-averse forecasts using quantile or super-quantile regression [Kirilenko2021]. In practice, FERC requires a quantile forecast value with a confidence level of 98% [FERC2024]. Recent work by [Kim2025] develops such quantile forecasts using a graph convolutional LSTM model.
@@
Despite these advances, existing models do not account for the actual costs of ETO nor provide a clear explanation on the effect of WP and DLR forecast error correlation. Without quantifying ETO-induced depreciation, models remain mathematically robust but economically incomplete. Previous works have handled the effect of ETO, and some have expanded it to DLR, but no works have integrated ETO costs into decision making such as OPF or UC. Harvey’s seminal work [Harvey1972] modeled the effect of ETO on aluminum conductors through an empirical equation for tensile strength loss (LoTS) as a function of conductor temperature and elapsed time. This model has been widely adopted to model ETO, including Musilek et al. [Musilek2012], EPRI [EPRI2005] and PJM [PJM2022]. [Musilek2012] analyzed the historical aging of a conductor using meteorological data. Moreover, [Teh2019] modeled the risk of DLR using the LoTS model and conducted a sensitivity analysis. However, this research does not consider the actual operation and performs only a simple calculation based on one line. A fretting-based reliability analysis model [Morozovska2020], points out the significance of corrosion modeling with ETO, but relies on large assumptions and limited empirical support. To the best of our knowledge, no existing optimization framework integrates realistic conductor health costs into grid operation.
@@
Incorporating ETO into optimization requires embedding a conductor temperature estimation model. However, owing to the nonlinear and multimodal (affected by many weather components) current-temperature relationship, direct integration into optimization is impractical. Previous studies have proposed conductor temperature models within OPF formulations [Kirilenko2021, Ngoko2018]. While these works employ current–temperature proxies, they do not capture the consequences of ETO, leaving a major gap in the literature.
@@
This paper addresses this gap by developing a day-ahead stochastic UC model that accounts for (i) DLR operation, (ii) correlated DLR and WP uncertainties, and (iii) the long-term costs of ETO. Incorporating ETO costs is challenging due to the state dependency of conductor depreciation models and the lack of a closed-form solution to the loading-temperature relationship. To overcome these challenges, we introduce a piecewise linear conductor depreciation cost function and a linear line loading–temperature proxy. The main contributions of this paper are as follows:
@@
    Modeling of ETO effects: We develop tractable models to capture the long-term effects of ETO, including (i) a robust and computationally efficient line loading–temperature proxy, and (ii) a risk-based piecewise conductor depreciation cost function.
@@
    Correlation risk characterization: We characterize how the correlation between WP and DLR forecast errors affects ETO-induced depreciation, providing theoretical justification for incorporating WP correlation in DLR risk-aware operation.
  4. 4.
    Validation on a realistic test system: We evaluate the CHA-UC on the Texas 123-bus backbone test system (TX-123BT) under uncertainty. Results highlight differences in commitment decisions and demonstrate improved handling of DLR and WP risks.
II Conductor cost depreciation model
This section quantifies the long-term effects of elevated conductor temperature operation (ETO) over the thermal limit and formulates a depreciation model suitable for optimization.
II-A Loss of tensile strength (LoTS) for ACSR lines
Long-term ETO causes conductor degradation, measured as the loss of tensile strength (LoTS). For aluminum strands, LoTS (%) is empirically modeled as shown in [Harvey1972]:
$L^{\mathrm{Al}}=\gamma(100-A\,t^{-\frac{25.4}{d}(0.001\tau-0.095)}),$ (1)
where $A=\min(100,\,134-0.24\tau)$, $\tau$ is conductor temperature (°C), $t$ is exposure time (h), and $d$ is conductor diameter (mm). $\gamma$, an environmental corrosivity factor, is introduced to reflect the effect of corrosive environments. For aluminum conductor steel-reinforced (ACSR) lines, the corresponding LoTS is approximated as given in [Musilek2012]: